1. A 30- ft board is cut into two parts such that the shorter piece is tree more than one half the longer.How long is each piece?
2. The perimeter of an isosceles triangle is 38 cm.The total length of the two equal legs is 6 more than thrice the length of the base.Find the length of three sides.
3. The width of a rectangle is one-third of the length. If the perimeter is 40 cm.,find the dimension of the rectangle.
thank u very much if u response....

Answer by rfer(10432)

1)
x+.5x+3=30
1.5x=27
x=18
.5x+3=12
2)
2(2x+6)+x=38
4x+12+x=38
5x=26
x=5.2
base=5.2
2 sides = 16.4 each
16.4+16.4+5.2=38
3)
2x+2(3x)=40
2x+6x=40
8x=40
x=5
proof
10+30=40